Can you please help me solve my problem?
{8*2-(3+9)]+[8-2*3]

Answer by checkley71(8403)

FIRST WE MULTIPLY ALL NUMBERS WITHIN THE ():
{8*2-(3+9)}+[8-2*3]
{16-(3+9)+[8-6]}
NOW WE ADD OR SUBTRACT THE NUMBERS WITHIN THE ()
{16-(12)+[2]} NOW WE ADD OR SUBTRACT THESE NUMBERS:
16-12+2=18-12=6 ANSWER.